• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E.Chudakov

    Log entry time 17:10:52 on June17,2005

    Entry number 146342

    keyword=Both septa tripped at 16:05

    At about 16:05 both septa magnets and the iocha49 tripped.
    The magnets quenched - the current went to zero and the counting
    rate in HRS dropped. The reason for the trip is not known.
    Rebooting the ioc (up to the point we could reset the magnets) took
    about 30 min. Now, we are ramping the magnets up at 6.A/min.





    FIGURE 1